AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  BRIKEN, Page 14 of 21  ELOG logo
  ID Date Author Subject
  154   Tue Nov 22 09:26:04 2016 A. Tarifeno-SaldiviaNoise check
Run for checking the noise just after turn on the crates, without high voltages on the PREAMPS.

Attached calibrated histogram for the full BRIKEN detector. 

Files:
161122_1820_000.dlt
161122_1820_000.root

Pulser counts: 1.179e4
  153   Tue Nov 22 08:28:36 2016 A. Tarifeno-Saldivia et alNew directory and Config File
A new directory in the IFICdaq computer has been created for the parasitic measurements during the Fallon Beam Time.

The path of the folder is:
/home/dacq/Desktop/Data/2211Fallon/

All the work during this campaign will be stored in folders inside here.

For the period before the parasitic experiments, the test and check of the detector will be stored in 
/home/dacq/Desktop/Data/2211Fallon/2211FallonPreparation

The Config file for the test is:

161107Conf_BrikenFull.xlsx
  152   Sun Nov 20 13:02:06 2016 Jorge, Alavaro, JoseDACQ
Gasific-7.0 Short Manual (Preliminary)
  151   Sun Nov 20 12:28:27 2016 Jorge, Alvaro, JoseDACQ
Updated run protocol
  150   Fri Nov 11 08:07:18 2016 BCR SG NTBBRIKEN Hit Visualization

These are the files used to create and visualize the hit patterns of the BRIKEN detector.

There are three files needed plus one input data file. The three files are "HybridBRIKENPhysicalHistogram.hpp", "HybridBRIKENPhysicalHistogram.cpp", and "HybryidBRIKEN_Tube_Info_1.txt". These files are attached to this eLog entry.

In order to run this type in a terminal

root -l HybidBRIKENPhysicalHistogram.cpp+

The plus is important, it tells ROOT to compile this with the gcc compiler and then run that compiled version of the code. It runs slow, if at all, in the ROOT C++ interpreter CINT.

The code has been tested with ROOT v5.32 and v5.34, not version 6.

At the bottom of the HybidBRIKENPhysicalHistogram.cpp file is the "main program". In it you can set the input data file, which should be a BRIKEN online root output file, and the 3He energy cuts. It is easiest if all the files are in the same folder. If there are any questions please feel free to email B. Charles (Charlie) Rasco.

  Draft   Thu Nov 10 05:05:54 2016 KMHow to convert ridf-file to root-file

1. Login to briken@ribfana03.riken.jp

2. Check if the ridf-file you want to convert is in the directory; /home/briken16/rawdata/bigrips/

3. Login to anaroot; type "anarootlogin"

4. Make your account as your name, and create your directory.

5. login to root; "root -l"

6.  root[ ] .L MakeFullBigRIPSTree_BRIKEN.C+

7.  root[ ]  MakeFullBigRIPSTree_BRIKEN("    full name of ridf-file    ",  Brho34 ,  Brho56 ,"  full name of output file  ")

You can check Brho34 and Brho56 in  --->  http://bripscnt01.rarfadv.riken.jp/runsum/index.php

 Find the link to the summary page in the "file" column. 

 For example: MakeFullBigRIPSTree_BRIKEN("radiioff0371.ridf",  7.176,   6.905, "test.root")

 

 

  148   Wed Nov 9 02:02:22 2016 IrisBoron rubber from India
The boron rubber from India is BR-2871, price US-$ 2700 for 2pc 1m x 2m x 3.3mm.
52% B content, density 1.61 g/cm3
  Draft   Wed Nov 9 01:04:35 2016 IrisDrawing of present setup
This is the setup which was used for the Ni runs (from Nov. 8th evening on).

Note (Nov. 10th): Alejandro built a U-shaped Pb wall, the degrader inside AIDA was removed and possibly a Al plate was installed. This has to be checked when access is possible again.


> Attached is the present setup including all walls, auxiliary detectors, degraders etc. and the Takechi setup.
> The lead wall will be removed for measurements (of course...).
> Not to scale! Some sizes and distances included. Pictures see Jose's entry ID 110.
> 
> Note the shape of the active area of the DeltaE Silicon detector, see previous entry.
  146   Wed Nov 9 01:02:53 2016 IrisElog
From Nov. 9th on we use the Takechi ELOG for the run information:
http://ribf-exp.riken.jp/elog/RIBF123/ 
  145   Tue Nov 8 17:00:27 2016 NTB, CG, ATSi Delta E

At least during the timeframe in between ~ 21:30 (11/08) - 23:30 (11/09) (161108_2323_717273Ni.root) we see that the silicon dE have a very low counting rate. ~1Hz when the first AIDA detector sees ~600 Hz and a cumulative ~2kHz. The energy range in the dE stops at ~10MeV.

Edit: It appears the Si dE were removed. Some signals are still in the spectra.

  144   Tue Nov 8 15:10:24 2016 Many717273Ni Run
F7 OUT F11 C-6mm

Start: 00:09
Run: 006
DLT: 161108_2251_717273Ni_006.dlt

Rates:
F11R: 1200 cps
BRIKEN: 270 cps (22.5%)

They  were tuning during this run.

Stop: 0042

Root file: 161109_0042_717273Ni.root
  143   Tue Nov 8 14:28:54 2016 Many717273Ni run
F7 CH2 F11 C-6mm
Total degrader thickness 6mm

Start:
RUN: 005
File: 161108_2251_717273Ni_005.dlt
Stop: 23:59
Run: 005

Rates:
F11R: 2600 cps
BRIKEN: 740 cps

Root file: 161109_0005_717273Ni.root


Counts: 
F11R: 5.36294e+06
BRIKEN: 1.42823e+06
Corr_BR_F11R: 1.41452e+06
  142   Tue Nov 8 13:49:58 2016 Many71,72,73Ni
F7 CH2 F11 Empty

We remove the 8mm Al plate from inside BRIKEN infront of AIDA
Move AIDA to the central position. We put a 5mm Al plate in front
of the degrader.

Degrader 5.4mm + 5mm

Start: 22:51
DLT: 161108_2251_717273Ni_03.dlt

We remove some degrader:2.2mm off
Then we start a new run: 004

Stop:23:26
To put the C-6mm target in F11

ROOT file: 161108_2323_717273Ni.root
  141   Tue Nov 8 12:51:08 2016 Many717273Ni
After removing the lead wall.

Start: 21:45
DLT: 161108_2145_717273Ni_001.dlt

During the first part we were changing the degrader.
Now they remove the C-6mm target.

So we increase the adjustable degrader to 5.4mm 
keeping the 8mm thick plate for implantation.

We stop and start a new DLT file at 22:02
DLT: 161108_2145_717273Ni_002.dlt
Stop: 22:48


NO ROOT FILE!
  140   Tue Nov 8 12:30:33 2016 Tain, Ale (who are not on shift)71,72,73Ni run
71,72,73Ni  run

We started at approximately 21:00, only online, not writing on tape. 

The lead is still in the way. We removed the degrader (fully) and the effect is a 5 % reduction of the neutron flux 
(determined using the rate of the rateF11/rateBRIKEN  before and after the change).
  139   Tue Nov 8 11:51:28 2016 Alespikes on some spectra
Some spectra show spikes or do not look normal. This should be checked.

Look at BRIKEN, Ring4, Ring5, Ring6, Ring7, D4, G7, Clovers, 
Rate1, Rate2, Rate3, etc.

Reset time stamp: 21.00, 8th-nov.

Probably there was a problem with the clock from AIDA related to the LN2 filling.
So to recover, we need to stop acq (only stop acquisition in DACQ control), 
send to hard (send parameters again) and once all the parameters were sent, 
Send Sync pulse.
  138   Tue Nov 8 08:49:51 2016 Tain et alMove ELOG
We will start to use the internal RIKEN ELOG
It is not clear that we can keep the present ELOG also updated
  137   Tue Nov 8 08:31:52 2016 Tain et al424344Ca run
Sequence of settings:
1) No target at F7 and Empty target at F11 (up to t=1E6ms)
2) No target at F7 and C-6mm target at F11 (up to t=5.1E6ms)
3) CH2 (10mm) target at F7 and C-6mm target at F11 (from t=5.3E6ms up to 6.8E6)
4) CH2 (10mm) target at F7 and Empty target at F11 (from t=6.8ms up to crash 7.1ms)

DAQ chash because start of Clover LN2 filling!
Counts accumulated in the the  same time channel at the end.

Start: 16:29 Run: 000
Stop: 18:28 Run: 000

DLT filename: 161108_1629_424344Ca_000.dlt
ROOT filename: 161108_1828_424344Ca.root
  136   Tue Nov 8 06:39:25 2016 Tain. IrisDrawing of present setup

Attached is the present setup including all walls, auxiliary detectors, degraders etc. and the Takechi setup.
The lead wall will be removed for measurements (of course...).
Not to scale! Some sizes and distances included. Pictures see Jose's entry ID 110.

Note the shape of the active area of the DeltaE Silicon detector, see previous entry.
  135   Tue Nov 8 06:13:11 2016 Tain et alNew plastic detector in F11
There is a new plastic scintillator
after the MUSIC chambers

The detector has a central hole of 108mm (with) x 38mm (height)
and is 3mm thick and is placed behind the existing plastic 
scintillator which is 5mm thick.
(see photo)
ELOG V3.1.4-unknown